# Preprocessing Disks for Convex Hulls, Revisited

### v1

@@ Abstract
In the preprocessing framework one is given a set of regions that one is allowed to preprocess to create some auxiliary structure such that when a realization of these regions is given, consisting of one point per region, this auxiliary structure can be used to reconstruct some desired output geometric structure more efficiently than would have been possible without preprocessing. Prior work showed that a set of $n$ unit disks of constant ply can be preprocessed in $O(n\log n)$ time such that the convex hull of any realization can be reconstructed in $O(n)$ time. (This prior work focused on triangulations and the convex hull was a byproduct.) In this work we show for the first time that we can reconstruct the convex hull in time proportional to the number of unstable disks, which may be sublinear, and that such a running time is the best possible. Here a disk is called stable if the combinatorial structure of the convex hull does not depend on the location of its realized point.
The main tool by which we achieve our results is by using a supersequence as the auxiliary structure constructed in the preprocessing phase, that is we output a supersequence of the disks such that the convex hull of any realization is a subsequence. One advantage of using a supersequence as the auxiliary structure is that it allows us to decouple the preprocessing phase from the reconstruction phase in a stronger sense than was possible in previous work, resulting in two separate algorithmic problems which may be independent interest. Finally, in the process of obtaining our results for convex hulls, we solve the corresponding problem of creating such supersequences for intervals in one dimension, yielding corresponding results for that case.
@@ 1 Introduction
The preprocessing framework for dealing with data uncertainty in geometric algorithms was initially proposed by Held and Mitchell [21]. In this framework, we have a set =$\mathcal{R}=\{R_{1},R_{2},\ldots,R_{n}\}$ of regions, often in $\mathbb{R}^{2}$, and a point set $P=\{p_{1},p_{2},\ldots,p_{n}\}$ with $p_{i}\in R_{i}$ (we also write $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$). This model has 2 consecutive phases: a preprocessing phase, followed by a reconstruction phase. In the preprocessing phase we have access only to $\mathcal{R}$ and we typically want to preprocess $\mathcal{R}$ in $O(n\log n)$ time to create some linear-size auxiliary data structure which we will denote by $\Xi$. In the reconstruction phase, we have access to $P$ and we want to construct a desired output structure $S(P)$ on $P$ using $\Xi$ faster than would be possible otherwise (see Figure 1). Löffler and Snoeyink [27] were the first to use this model as a way to deal with data uncertainty: one may interpret the regions $\mathcal{R}$ as imprecise points, and the points in $P$ as their true (initially unknown) locations. This interpretation of the preprocessing framework has since been successfully applied to various problems in computational geometry [7, 10, 17, 26, 41, 40].
@@
The convex hull of a set of points in $\mathbb{R}^{2}$ is the smallest convex set that contains all points. The algorithmic problem of computing the convex hull is arguably one of the most fundamental problems in computational geometry [2], and has been considered in an impressive range of different models for geometric uncertainty [19, 20, 30, 13, 22, 28, 16, 17, 24, 11, 3, 23]. The first result for convex hulls within the preprocessing framework can in fact be derived from the work by Held and Mitchell [21], who show that a set of disjoint unit disks can be preprocessed in $O(n\log n)$ time such that a triangulation can be reconstructed in linear time: since any triangulation contains at least the edges of the convex hull the convex hull can be extracted in the same amount of time (see Figure 1). Similarly, the work by van Kreveld et al. [41] implies that a set of disjoint disks of arbitrary radii (or a set of disjoint polygons of constant complexity) can also be preprocessed in $O(n\log n)$ time for the same purpose, and the work by Buchin et al. [6] implies that the same is true for a set of moderately overlapping unit disks. While these results focus on different problems (triangulation or Delaunay triangulation) and the convex hull is “only a by-product”, Ezra and Mulzer [17] explicitly study the convex hull and show that any set of lines in the plane can also be preprocessed to speed up the reconstruction of the convex hull, although not to linear time: they achieve a reconstruction time of $O(n\alpha(n)\log^{*}n)$ in expectation.
@@
Motivated by these observations, van der Hoog et al. [39, 40] have recently explored how the preprocessing model may be modified to allow for such sublinear reconstruction; for instance, for sorting they achieve reconstruction time proportional to the entropy of the interval graph, which can be less than $n$ if many intervals are isolated. In order to support sublinear reconstruction, van der Hoog et al.introduce an additional phase to the preprocessing framework, see Section 4.2 for more details.
Figure 2: (a) A set of disjoint unit disks $\mathcal{R}$. (b) All possible combinatorial convex hulls. We identify five types of disks: (i) stable impossible interior disks that never contribute to the convex hull (red); (ii) unstable potential interior disks that may or may not contribute (orange); (ii) unstable potential boundary disks that may or may not contribute (yellow); (iii) unstable guaranteed boundary disks that are guaranteed to contribute, but for which the location may influence the structure of the hull (blue); and (iv) stable guaranteed boundary disks that are guaranteed to contribute, and the structure is independent of their location (green). (Refer to Section 2.1 for precise definitions).
1.1 Our contribution
Preprocessing disks for convex hulls.
In this work, we investigate the problem of preprocessing a set of disks in the plane to reconstruct the convex hull directly for the first time. Doing so allows us to achieve features that are not possible when viewing the convex hull as a by-product of a triangulation, such as potential sub-linear reconstruction times. For some sets of disks we can already identify a subset of disks that is guaranteed to not take part in the convex hull, and a subset of disks that is guaranteed to take part in the convex hull (see Figure 2), and thus we need not spend time on them during the reconstruction phase.
Figure 3: (a) A set of disjoint unit disks $\mathcal{R}$ and (b) a sequence of disks that is guaranteed to contain the vertices of the convex hull in the correct order. (c) A possible true convex hull.
@@ 1.1 Our contribution
As the main tool in our solution, we explore a new generic auxiliary structure: the supersequence. In principle, it is applicable to any computational problem where the output structure $S(P)$ is an ordered subset of $P$; that is, $S(P)=\langle p_{i_{1}},p_{i_{2}},\ldots,p_{i_{s}}\rangle$ for some indices $\{i_{1},i_{2},\ldots,i_{s}\}\subseteq[n]$. For such problems, an attractive option to use for the auxiliary structure $\Xi$ is a supersequence of the regions corresponding to $S$; that is, $\Xi$ is a sequence of (possibly reccuring) elements of $\mathcal{R}$ with the guarantee that, no matter where the true points $P$ lie in their regions, the sequence of elements of $P$ which we would obtain by replacing the regions in $\mathcal{R}$ by their points will always contain $S(P)$ as a subsequence. In the case of the convex hull $S(P)$ is the vertices of the convex hull in counter-clockwise order (the fact that this is a cyclic order slightly complicates things; in our solution we instead compute four quarter hulls separately) and $\Xi(P)$ is a sequence of disks (see Figure 3).
@@
An advantage of using a supersequence as auxiliary structure is that it allows us to decouple the preprocessing phase from the reconstruction phase into two separate algorithmic problems of independent value, in a stronger sense than was possible in previous work. In heriditary algorithms, introduced by Chazelle and Mulzer [8], one is given a structure on a superset, and is interested in computing the structure on a subset. Heriditary algorithms have been used in the context of the preprocessing framework, for instance, van Kreveld et al. [41] in their reconstruction phase first produce a triangulation of red and blue points, where the red points are the points of interest, and then describe a linear-time hereditary algorithm for triangulations to obtain the final result. Our approach of using supersequences is different in that we do not require explicit construction of a structure on a superset; instead, in the reconstruction phase we can trivially replace the disks by their points and are left with a pure computational problem (Open Problem 2.15, in the case of convex hulls).
Preprocessing intervals for sorting.
As a subroutine in our solution, we require to preprocess certain sets of moderately overlapping unit intervals into a supersequence such that when we get a single value per interval, the sorted order can be reconstructed quickly. We believe this result is also of independent interest; refer to Section 3 for more background on this problem.
@@ 1.2 Organization
  • In Section 2, we first establish the necessary vocabulary to be able to state our main results, which includes a classification of the input disks that is interesting in its own right. We then state our results for preprocessing unit disks for the convex hull, and discuss some of their limitations.
  • Before proceeding with the proofs of our main results, in Section 3, we first switch to the one-dimensional problem of preprocessing a set of intervals for sorting, which is required as a subroutine later, but is also of independent interest. Our main result here is an algorithm to produce a supersequence of intervals that is guaranteed to contain any instance in sorted order.
  • In Section 4, we then consider the corresponding problem of reconstructing the sorted order from such a sequence. In particular, we show that it is always possible to recover the sorted order from a smooth supersequence in linear time, but for general sequences this depends on the model of computation.
  • Then, in Section 5 we return to the two-dimensional problem and discuss how to preprocess a set of disjoint unit disks as a supersequence of the convex hull. In Section 6 we extend this to overlapping unit disks using the one-dimensional result.
  • In Section 7, we consider the corresponding reconstruction problem. Again, we show that it is always possible to recover the convex hull from a smooth supersequence. Whether it is possible for general supersequences is left as an open problem.
  • Finally, in Section 8, we summarize our findings and discuss directions for future research.
2 Preliminaries, Results, and Limitations
In this section we formally define the problem and state our result for preprocessing unit disks of bounded ply in $\mathbb{R}^{2}$ for reconstructing the convex hull. Before we can state our results, we need to first discuss a classification of the input disks into five types, a notion of quarter hull to decompose the problem into smaller parts, and a notion of smoothness.
2.1 Disk Characterization
Figure 4: The same example as Figure 2 with the regions $I$ and $E$ drawn in.
We may classify disks [^1] Although we mostly consider unit disks in this work, the classification in this section applies to (potentially overlapping) disks of any radii. in the plane in multiple ways with respect to the convex hull in the preprocessing framework. First, we classify the disks on whether their corresponding points will appear on the convex hull or not. We will define impossible, potential, and guaranteed disks.
@@ 2.1 Disk Characterization
Prior work has considered the union or intersection of all possible realizations of $\mathcal{R}$ under different names (see for example [29, 33]), and these structures are closely related to our above classification of disks. In particular, the above definitions can be equivalently phrased in terms of inclusion or exclusion in certain regions; the following observations have been made before. Let $I$ be the intersection of all halfplanes that intersect all disks. Let $E$ be the intersection of all halfplanes that fully contain at least $n-1$ disks. See Figure 4. {observation} A disk is impossible if and only if it is contained in $I$, and a disk is guaranteed if and only if it lies outside $E$.
@@
A disk $D\in\mathcal{R}$ is called stable, if for any fixed realization $P^{\prime}$ of =$\mathcal{R}^{\prime}=\mathcal{R}\setminus\{D\}$, the combinatorial structure [^2] The ordered list of disks corresponding to the clockwise sorted order of the vertices of the convex hull. of the convex hull of $P^{\prime}\cup\{p\}$ is the same for any realization $p$ of $D$. A disk that is not stable is unstable.
@@
These 3 different classifiers for each disk, in principle give $3\times 2\times 2=12$ different types of disks. However, due to certain dependencies many of these types are not possible.
Lemma 2.4.

### v2

@@ Abstract
In the preprocessing framework for dealing with data uncertainty, one is given a set of regions that each represent one uncertain data point. Then, one is allowed to preprocess these regions to create some auxiliary structure, such that when a realization of the regions is given, this auxiliary structure can be used to reconstruct some desired output geometric structure more efficiently than would have been possible without it. A classic result in this framework is that a set of $n$ unit disks of constant ply can be preprocessed in $O(n\log n)$ time such that the convex hull of any realization can be reconstructed in $O(n)$ time. In this work we revisit this problem and show that, in fact, we can reconstruct the convex hull in time proportional to the number of unstable disks, which may be sublinear, and that such a running time is the best possible. A disk is stable if the combinatorial structure of the convex hull does not depend on the location of its realized point. We also extend our results to overlapping disks, and disks of varying radii.
The main tool by which we achieve our results is by using a supersequence as the auxiliary structure constructed in the preprocessing phase. That is, we output a sequence of the disks (possibly with repetitions) such that the convex hull of any realization is guaranteed to be a subsequence of this sequence. Using a supersequence as the auxiliary structure allows us to decouple the preprocessing phase from the reconstruction phase in a stronger sense than was possible in previous work, resulting in two separate algorithmic problems which may be independent interest.
@@ 1 Introduction
The preprocessing framework for dealing with data uncertainty in geometric algorithms was initially proposed by Held and Mitchell [22]. In this framework, we have a set $\mathcal{R}=\{R_{1},R_{2},\ldots,R_{n}\}$ of regions, often in $\mathbb{R}^{2}$, and a point set $P=\{p_{1},p_{2},\ldots,p_{n}\}$ with $p_{i}\in R_{i}$ (we also write $P\mathrel{\subset}\mkern-10.0mu\joinrel\sim\mathcal{R}$). This model has 2 consecutive phases: a preprocessing phase, followed by a reconstruction phase. In the preprocessing phase we have access only to $\mathcal{R}$ and we typically want to preprocess $\mathcal{R}$ in $O(n\log n)$ time to create some linear-size auxiliary data structure which we will denote by $\Xi$. In the reconstruction phase, we have access to $P$ and we want to construct a desired output structure $S(P)$ on $P$ using $\Xi$ faster than would be possible otherwise (see Figure 1). Löffler and Snoeyink [29] were the first to use this model as a way to deal with data uncertainty: one may interpret the regions $\mathcal{R}$ as imprecise points, and the points in $P$ as their true (initially unknown) locations. This interpretation of the preprocessing framework has since been successfully applied to various problems in computational geometry [7, 10, 16, 28, 48, 47].
@@
The convex hull of a set of points in $\mathbb{R}^{2}$ is the smallest convex set that contains all points. The algorithmic problem of computing the convex hull is arguably one of the most fundamental problems in computational geometry [38], and has been considered in an impressive range of different models for geometric uncertainty [18, 20, 37, 13, 23, 31, 15, 16, 25, 11, 1, 24] (refer to Appendix A). The first result for convex hulls within the preprocessing framework can in fact be derived from the work by Held and Mitchell [22], who show that a set of disjoint unit disks can be preprocessed in $O(n\log n)$ time such that a triangulation can be reconstructed in linear time: since any triangulation contains at least the edges of the convex hull, the convex hull can be extracted in the same amount of time (see Figure 1). Similarly, the work by van Kreveld et al. [48] implies that a set of disjoint disks of arbitrary radii (or a set of disjoint polygons of constant complexity) can also be preprocessed in $O(n\log n)$ time for the same purpose, and the work by Buchin et al. [6] implies that the same is true for a set of moderately overlapping unit disks. While these results focus on different problems (triangulation or Delaunay triangulation) and the convex hull is “only a by-product”, Ezra and Mulzer [16] explicitly study the convex hull and show that any set of lines in the plane can also be preprocessed to speed up the reconstruction of the convex hull, although not to linear time: they achieve a reconstruction time of $O(n\alpha(n)\log^{*}n)$ in expectation.
@@
Motivated by these observations, van der Hoog et al. [46, 47] explore how the preprocessing model may be modified to allow for such sublinear reconstruction; for instance, for sorting they achieve reconstruction time proportional to the entropy of the interval graph, which can be less than $n$ if many intervals are isolated. In order to support sublinear reconstruction, van der Hoog et al.introduce an additional phase to the preprocessing framework, see Section 5.2 for more details.
Figure 2: (a) A set of disjoint unit disks $\mathcal{R}$. (b) All possible combinatorial convex hulls. We identify five types of disks: (i) stable impossible interior disks that never contribute to the convex hull (red); (ii) unstable potential interior disks that may or may not contribute (orange); (ii) unstable potential boundary disks that may or may not contribute (yellow); (iii) unstable guaranteed boundary disks that are guaranteed to contribute, but for which the location may influence the structure of the hull (blue); and (iv) stable guaranteed boundary disks that are guaranteed to contribute, and the structure is independent of their location (green). (Refer to Section 2 for precise definitions).
1.1 Contribution
@@
In previous work in the preprocessing framework, the auxiliary structure, which is produced during the preprocessing phase, can be anything; and indeed very different techniques and have been used successfully in the past. While this illustrates the breath of the framework, it also makes it hard or impossible to separate the preprocessing phase from the reconstruction phase: they only make sense when viewed as a whole.
@@
In this work, we introduce a new generic auxiliary structure: the supersequence. In principle, it is applicable to any computational problem where the output structure $S(P)$ is an ordered subset of $P$; that is, $S(P)=\langle p_{i_{1}},p_{i_{2}},\ldots,p_{i_{s}}\rangle$ for some indices $\{i_{1},i_{2},\ldots,i_{s}\}\subseteq[n]$. For such problems, an attractive option to use for the auxiliary structure $\Xi$ is a supersequence of the regions corresponding to $S$; that is, $\Xi$ is a sequence of (possibly reccuring) elements of $\mathcal{R}$ with the guarantee that, no matter where the true points $P$ lie in their regions, the sequence of elements of $P$ which we would obtain by replacing the regions in $\mathcal{R}$ by their points will always contain $S(P)$ as a subsequence.
A supersequence is arguably a more meaningful structure than the structures used in previous work; as a result, algorithms for the two phases may also be of independent interest, outside of the preprocessing framework. In addition, the supersequence viewpoint naturally allows us to incorporate sublinear reconstruction times in certain settings, something that was not possible in the earlier work and has only been explored more recently. We believe the concept is potentially applicable to many other geometric problems (for example, the Pareto front, the TSP tour, or the shortest path in a polygon).
@@
This idea is somewhat related to the concept of heriditary algorithms, introduced by Chazelle and Mulzer [8], in which one is given a structure on a superset, and is interested in computing the structure on a subset. Heriditary algorithms have been used in the context of the preprocessing framework, for instance, van Kreveld et al. [48] in their reconstruction phase first produce a triangulation of red and blue points, where the red points are the points of interest, and then describe a linear-time hereditary algorithm for triangulations to obtain the final result. Our approach of using supersequences is different in that we do not require explicit construction of a structure on a superset; instead, in the reconstruction phase we can trivially replace the disks by their points and are left with a pure computational problem (Open Problem 1.2, in the case of convex hulls).
Convex hull supersequences.
In this work, we focus on the convex hull as our structure of interest; that is, $S(P)$ is the sequence of vertices of the convex hull, [^1] For technical reasons we actually compute four quarter hulls separately; see Section 3.1 and $\Xi(P)$ is a sequence of disks (see Figure 3). It is worth noting that the notion of a supersequence is flexible enough to allow some disks to appear in the sequence very often, or even a linear number of times (which in fact would be necessary when dealing with arbitrary disks rather than unit disks), as long as the total sequence length is not too large.
In order to support sublinear reconstruction, we must identify subsets of disks that are guaranteed to not take part in the convex hull, and subsets of disks that are guaranteed to take part in the convex hull in a particular order (see Figure 2). In Section 2 we formalize these notions and discuss how they relate to other properties of set of disks that have been used in the past, as well as discuss computational aspects.
We study several variants of the problem: input regions are disjoint unit disks, partially overlapping disks, or disks of varying sizes. Our results are formally stated in Section 3.3.
Figure 3: (a) A set of disjoint unit disks $\mathcal{R}$ and a sequence of disks (indices) that is guaranteed to contain the vertices of the convex hull in the correct order. (b) A possible true convex hull, and the corresponding subsequence. (c-d) An example with disjoint disks of different radii; note that $R_{6}$ must appear in the sequence a linear number of times.
1.2 Limitations & Open Problem
We now discuss some limitations and an open problem. Note that we assume we are working in the real RAM model of computation, which is consistent with earlier work in our framework.
Restrictions on the supersequences.
[…]
Note that, although we require handling duplicates in our application, Problem 1.2 is also open even when all points are unique. We believe this question is of independent interest.
General position.
For ease of exposition, throughout we assume general position of the disks in the preprocessing phase and general position of the true points in the reconstruction phase. As mentioned in Section 2, general position for disks means that disks have non-zero radius, no two disk centers coincide, and no three disks are tangent to a common line. For points, general position means all points are distinct and no three points are collinear.
However, because we will actively duplicate regions and points in our sequences, we will need to drop (part of) this assumption in some subroutines. We expect our general position assumptions may be completely lifted by using symbolic perturbation [14]; however, this will certainly complicate the discussion by introducing additional cases in several arguments.
@@ 1.3 Organization
In Section 2 we first discuss how uncertainty regions modeled as disks can be classified into types based on their potential roles in the convex hull, which is of independent interest outside the context of the preprocessing framework.
Then in Section 3 we formally define the problem and state our result for preprocessing disks of bounded ply and radii in $\mathbb{R}^{2}$ into a supersequence which can be used for efficiently reconstructing the convex hull.
In Section 4, we discuss how to preprocess our set of disks. We first discuss some preliminaries, and then consider disjoint unit disks, overlapping unit disks, and finally disks of varying (though bounded) radii.
In Section 5 we discuss the reconstruction phase. First, in Section 5.1 we show that it is always possible to recover the convex hull from a smooth supersequence. Whether it is possible for general supersequences is left as an open problem. Finally, in Section 5.2 we show how to use our classification from Section 2 to reconstruct the convex hull in time proportional to the number of unstable disks.
In Section 6, we summarize our findings and discuss directions for future research.
2 Classification of Uncertainty Disks for Convex Hulls
In this section we discuss a classification of uncertainty regions modeled as disks into five types, depending on their potential roles in the convex hull. The results in this section are in principle independent of the preprocessing framework.
Throughout, for a subset $X$ of objects in the plane (i.e. regions or points), we write $\mathrm{CH}(X)$ to denote the convex hull of $X$.
Figure 4: A set of disks and their classification, and the regions $I$ and $E$.
We may classify disks in the plane in multiple ways with respect to the convex hull in the preprocessing framework. For ease of description, we will assume general position: all disks have positive (nonzero) radius, no two disks have a common center, and no line is co-tangent to three or more disks or disk centers.
First, following e.g. [5], we classify the disks on whether their corresponding points will appear on the convex hull or not. We will define impossible, potential, and guaranteed disks.
@@ 2 Classification of Uncertainty Disks for Convex Hulls
Prior work has considered the union or intersection of all possible realizations of $\mathcal{R}$ under different names (see for example [36, 42]), and these structures are closely related to our above classification of disks. In particular, the above definitions can be equivalently phrased in terms of inclusion or exclusion in certain regions; the following observations have been made before. Let $I$ be the intersection of all halfplanes that intersect all disks. Let $E$ be the intersection of all halfplanes that fully contain at least n1n-1 disks. See Figure 4. {observation} A disk is impossible if and only if it is contained in $I$, and a disk is guaranteed if and only if it lies outside $E$.
@@
A disk $D\in\mathcal{R}$ is called stable, if for any fixed realization $P^{\prime}$ of $\mathcal{R}^{\prime}=\mathcal{R}\setminus\{D\}$, the combinatorial structure [^2] The ordered list of disks corresponding to the clockwise sorted order of the vertices of the convex hull. of the convex hull of $P^{\prime}\cup\{p\}$ is the same for any realization $p$ of $D$. A disk that is not stable is unstable.
@@
[…]
Let $\mathcal{B}(\mathcal{R})=\langle B_{1},\ldots,B_{k}\rangle$ be the sequence of boundary disks, indexed in counterclockwise order of appearance on the convex hull, see Figure 5(a). Note that not every disk in $\mathcal{R}$ needs to appear in $\mathcal{B}(\mathcal{R})$, and furthermore, for non-unit disks, the same disk may appear multiple times, that is we may have $B_{i}=B_{j}$ for $i\neq j$. It is known that $k\leq 2n-1$ [40]. For any $i\in[1,k]$, define the $i$th strip as $s_{i}=\mathrm{CH}(B_{i},B_{i+1})$ (where $s_{k}=\mathrm{CH}(B_{k},B_{1})$). We refer to the chain of strips, ordered by increasing $i$, as the strip hull of $\mathcal{R}$. (Indeed the convex hull of all strips is equal to that of all regions, where straight segments on the boundary follow individual strips and turns occur at the intersection of the common disk with the next strip.) Furthermore, for any $i$, we call the line segment connecting the centers of $B_{i}$ and $B_{i+1}$ the spine of the corresponding strip. Observe that the spines themselves define a plane geometric graph, which we call the spine hull, see Figure 5(c). Note that when $\mathcal{R}$ is a set of unit disks, the spine hull is simply the convex hull of the disk centers.
The classifications in Definitions 2.1, 2.2, and 2.3, in principle give $3\times 2\times 2=12$ different types of disks. However, due to certain dependencies many of these types are not possible.
{observation}
@@
Proof 2.4.
First, observe that a guaranteed disk $D\in\mathcal{R}$ must be a boundary disk, as by Observation 2, $D$ must lie outside $E$, or equivalently outside the convex hull of $\mathcal{R}\setminus\{D\}$.
Second, a potential disk $D\in\mathcal{R}$ must be unstable, as $D$ being potential means there is both a realization of $\mathcal{R}$ where $D$ is realized as a vertex of the convex hull and a realization of $\mathcal{R}$ where it is not a vertex of the convex hull, and thus the combinatorial structure depends on the realization of DD. [^3] The definition of stable does not directly prohibit a scenario where for some fixed realization of $\mathcal{R}\setminus\{D\}$, any realization of $D$ is always a vertex of the convex hull, while for some other fixed realization of $\mathcal{R}\setminus\{D\}$, any realization of $D$ is never a vertex of the convex hull. However, such a scenario is impossible due to the continuity of our disk based uncertain regions.
Finally, an impossible disk must be both interior and stable. Clearly it is stable as its realization is never a vertex of the convex hull and thus does not affect its combinatorial structure. It is interior as by Observation 2 it lies inside $I$ which lies strictly inside the conex hull of $\mathcal{R}$.
Observation 2 implies there are only 5 possible types of disks (see Figure 4).
2.1 Computation of the Disk Classification
Here we argue that the disk classifications into the 5 possible types from Section 2 can be computed in $O(n\log n)$ time. First, we will narrow down the possible configurations of disks in which stable guaranteed boundary disks can appear.
{observation}
A guaranteed disk can appear at most once on the strip hull.